# Normal Quaternionic Matrices and Finitely Generated Witt Rings

### v1

@@ 5. Normal Quaternionic Matrices
Let $M=(M_{i,j})_{0\leq i,j\leq n-1}$ be a normal quaternionic matrix of type $\epsilon\in\{0,1\}$ with quaternionic structure $S:=S(M):=(G,-1,q)$. Set $x:=\dim(V(-1))$ if $-1\not=1\in G$ and otherwise $x:=\operatorname{max}\{\dim(V(a))\mid a\in G\setminus\{1\}\}$. We write $s:=s(S)$. Then
@@
The values $(q(a_{0},a_{j})\mid j\in J)$ must be independent for else there would be a subset $I\subset J$ such that
@@
Proposition 5.1 directly produces a list of ’potential’ normal quaternionic matrices of order $1$ and $2$ and the different matrices are clearly not isomorphic as the dimension of the value sets $V(a)$ is invariant under isomorphism. For type 00, we get exactly the matrices
@@
Given $x,y\in\mathbb{N}$ we say that $y$ is a factor of $x$ if $x*y<x$.
@@
By Proposition 5.1, the first row of $M$ has the asserted shape, setting $d:=n-x$ if $s\leq 2$ and $d:=n-x-1$ otherwise.
We write $S:=S(M)=(G,-1,q)$ and fix a basis $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ of $S$ such that $M_{\mathcal{A}}(S)=M$. To keep notation brief, we usually only write $q$ for the quaternionic map instead of $q_{\mathcal{B}}$ for a basis $\mathcal{B}$ of the full 2B-group $B$ group whenever no confusion is imminent.
@@
Permuting the $u-t$ columns in front of the last $t-r$ columns of $B_{4}$, we can thus attain a block of the shape
@@
(c): By Proposition 5.1, the dimension $\dim(V(a_{0}))=n-d$ is maximal. If $d\not=n-1$, it follows that $M_{0,0}=M_{1,0}=M_{0,1}=M_{1,1}=0$ and counting the number of zeroes in the second row with Proposition 5.5, we get $\dim(V(a_{1}))=n-r-v+t$. We thus have $n-r-v+t\leq n-d$ and the claim follows. ∎
@@
This can be seen as the rows with index $3$ resp. $5$ of $\overline{M}$ are (0 1 0 1 2 3 2 3)01012323(0\ 1\ 0\ 1\ 2\ 3\ 2\ 3)( 0 1 0 1 2 3 2 3 ) and (0 2 1 3 0 2 1 3)02130213(0\ 2\ 1\ 3\ 0\ 2\ 1\ 3)( 0 2 1 3 0 2 1 3 ), so the entries $1$ and $2$ appear in both rows but do not appear in common columns.
@@
    For direct products, the situation is more involved. Take structures $S=(G,-1,q)$ and $S^{\prime}=(G^{\prime},-1,q^{\prime})$ with normal quaternionic matrices $M=M_{\mathcal{A}}(S)$ and $M^{\prime}=M_{\mathcal{A}^{\prime}}(S^{\prime})$ for bases $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ of $S$ and $\mathcal{A}^{\prime}=(a_{0}^{\prime},\ldots,a_{n^{\prime}-1}^{\prime})$ of $S^{\prime}$. We conjecture that we get $M(S\times S^{\prime})=M_{\mathcal{C}}(S\times S^{\prime})$ for a basis $\mathcal{C}=(c_{0},\ldots,c_{n+n^{\prime}-1})$ that is close to a union of $\mathcal{A}$ and $\mathcal{A^{\prime}}$ in the following way: Consider $G$ and $G^{\prime}$ as subsets of $G\times G^{\prime}$ canonically. Then $c_{0}\in\{a_{0},a_{0}^{\prime},a_{0}a_{0}^{\prime}\}$ depending on the levels of $S$ and $S^{\prime}$ and
[…]
@@
The situation for direct products also gives rise to quaternionic matrices of elementary type structures where nonzero entries appear that are no powers of 2. This is due to the fact that we have to fix $-1$ as the first basis element. For example, we get
@@ 6. The Computation
We start with the following reduction technique. Once the ETC has been confirmed for order up to $n-1$, in order to verify the ETC for order $n$, it is enough to show the ETC for nondegenerate structures that are neither rigid nor pythagorean by Proposition 2.8(a) and (d). Furthermore, depending on the level, we can also exclude some $2$-rigid cases. To make use of this, we include the following results that show how how to read these properties off normal quaternionic matrices.
@@
Let $S$ be a quaternionic structure of order $n$ with normal quaternionic matrix $M$. Then $S$ is nondegenerate and neither rigid nor pythagorean if and only if every row of $\overline{M}$ has at least $4$ and at most $2^{n-1}$ zeroes.
@@
The first part follows immediately from the definitions as the zeroes in a row of $\overline{M}$ correspond $1-1$ to the elements of $V(a)$ for some $a$. Considering Proposition 5.1 along with (b)(M2) and (b)(M3), we get the second part of the assertion. ∎
@@
Let $S=(G,-1,q)$ be a nondegenerate quaternionic structure of order $n$ with normal quaternionic matrix $M$. We have $s(S)=1$ if and only if $M$ has only zeroes on the diagonal. Else, we have $s(S)=2$ if and only if $M_{0,0}=0$. Lastly, we get $s(S)=4$ if and only if $M_{0,0}=1=\overline{M}_{I,J}$ for some $I,J\subset\{1,\ldots,n-d-2\}$ where $d$ is defined in the same way as in Proposition 5.5, i.e. $\overline{M}_{\{0\},I}=\overline{M}_{\{0\},J}=0$.
@@
$-1\in D(\langle 1,1\rangle)\iff q(-1,-1)=0\iff M_{0,0}=0.$
If $s(S)=4$, we have $M_{0,0}=1$ as seen before and $-1\in D(\langle 1,1,1,1\rangle)$. Note that by the correspondence of $q$ and $M$ (cf. 4.4) the existence of $I,J$ as asserted above is equivalent to the existence of $a,b\in G$ such that $q(a,b)=1$ and $q(-1,a)=q(-1,b)=0$. By a classical property which also holds in the abstract setting (cf. [Marshall_AbstractWittRings, Proposition 2.10]), we have
@@
Hence, we find $a\in D(\langle 1,1,1\rangle)$ and $b\in D(\langle 1,1\rangle)$ such that $q(-a,-1)=q(-b,a)=q(-1,b)=0$. By bimultiplicativity and symmetry of $q$, and applying $q(-1,-1)=1$, it follows that $q(-1,a)=1$ and $q(a,b)=q(a,-1)\ast q(a,-b)=1$. As pointed out above, the existence of $a,b$ proves the existence of $I,J$ as asserted. Conversely, if there are $a,b\in G$ such that $q(a,b)=1$ and $q(-1,a)=q(-1,b)=0$, we can do the same computations as above to show that $-1\in D(\langle 1,a\rangle)$ and $a\in D(\langle 1,1,1\rangle)$, completing the proof. ∎
@@
The previous two results enable us to apply the results stated in Proposition 2.8 to exclude some cases. We explicitly collect these here for a better overview. To confirm the ETC for order $n$, whenever we know that the ETC holds for order up to $n-1$, we can exclude quaternionic structures whose normal quaternionic matrix $M$ has one of the following properties:
  • There is a row of the completion $\overline{M}$ that has only zeroes or fewer than $4$ zeroes.
  • We have $M_{n-1,n-1}=0$ (i.e. level $1$) and any row of the completion has fewer than $8$ zeroes.
  • We have $M_{0,0}$ (i.e. level at most $2$) and any row of the completion starting with 00 has fewer than $8$ zeroes.
  • We have $M_{0,0}=M_{1,2}=1$ and $M_{0,3}=2$ or $n=3$. This means that $-1$ is $2$-rigid and we have level $4$, corresponding to the parameters (cf. Proposition 5.5) $d=n-3$, $f=1$, $r=1$ arbitrary $v,t,u$. Note that in this case, we always have $v=n-2$ by the restrictions presented in Corollary 5.6.
@@
  1. (a)
    All quaternionic structures of order up to $6$ are of elementary type.
  2. (b)
    All quaternionic structures of order $7$ and level in $\{1,2\}$ are of elementary type. Furthermore all those of order $7$ and level at least $4$ with parameters (cf. Proposition 5.5)
    $(d,f,r,v,t,u)\not\in\Big{\{}(3,1,0,3,0,0),\quad(4,1,0,4,1,1),\quad(4,1,0,4,2,2),\quad(4,2,1,4,1,1)\Big{\}}$
    are of elementary type.
@@
(a): By the results of Carson and Marshall as cited in 2.8(c), every quaternionic structure of order up to $5$ is of elementary type. Hence, it suffices to compute the correct upper bound for the nonpythagorean, nonrigid and nondegenerate normal quaternionic matrices, also excluding the existence of $2$-rigid elements in the cases mentioned in Lemma 2.9. By Remark 2.10 there are $81=\sum_{k>1}e_{k}(6)$ nonrigid and nondegenerate normal quaternionic matrices of order $6$. One readily checks that among those, there are $10$ of level $1$, $28$ of level $2$ and $43$ of level at least $4$.
@@
Applying this, we verified the ETC for order $6$ by running Algorithm 1 with the above restrictions, computing that there are at most $4$ resp. $19$ resp. $43$ normal quaternionic matrices of order $6$ and level $1$ resp. $2$ resp. at least $4$.
The argument for (b) is the same, building further on the results of part (a). Here, the lower bound on the number of normal quaternionic matrices given by the elementary type structures is $4$ resp. $65$ resp. $199$ for level $1$ resp. $2$ resp. at least $4$. For level $1$ and $2$, these numbers are obtained as upper bounds by Algorithm 1. For level $4$, excluding the parameters
$(3,1,0,3,0,0),\quad(4,1,0,4,1,1),\quad(4,1,0,4,2,2),\quad(4,2,1,4,1,1),$
we obtain $198$ normal quaternionic matrices which verifies the ETC in these cases since there is one elementary type structure with parameters $(3,1,0,3,0,0)$, namely $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta_{3}$, and no elementary type structures for the other excluded parameters. ∎
We briefly outline some pseudo code in Algorithm 1 to classify normal quaternion matrices while the complete version we used can be found via the following link: https://github.com/nicolo314/NormalQuaternionicMatrices.
In the program, partial quaternionic matrices (or what is aimed to be completed to such) are created row by row and via depth search in order to minimize the data that has to be stored simultaneously. To keep track of the current object, we save our current data in an $n\times n$ matrix $M$ whose entries are tuples. The first entry of the tuple in $M(i,j)$ is a partial quaternionic matrix $A$ with entries up to position $(i,j)$ and the second entry of $M(i,j)$ is a list of possible extensions at index $(i,j+1)$ of $A$ that still have to be handled. We will make use of functions that increase resp. decrease $(i,j)$ to the next resp. last index, taking into considerations that elements of a quaternionic matrix up to the diagonal of a given row are already determined by previous entries.
@@
              if Validity Check($A$) = true  then
                  $A\leftarrow$Add new row($A$)
@@
                  $M(i,j)\leftarrow(A,\,$possible next elements($A$)))))
@@
                       else if $j=n-1$ then
                           if $i=n-1$ then
[…]
@@
                                $A\leftarrow$Add new row($A$)
                                $M(i,j)\leftarrow(A,\,$possible next elements($A$)))))
@@
         return $A$ appended by partial row $k+1$ up to diagonal using (b)(M2) and (b)(M3)
function possible next elements($A$)
     if Validity Check($A$) = true  then
@@
  1. (a)
    In order to tap the full potential of a modern computer architecture with multiple cores, we parallelized some of our calculations. Mostly, we simply parallelized the different choices for parameters $s,d,f,r,t,u,v$, i.e. the different choices for the first two rows, as presented in Proposition 5.5. This works well if there are no parameters with tremendously large runtime, so in particular for $n\leq 6$.
  2. (b)
    For $n=7$ and certain sets of parameters, we used a more involved strategy that sped up the run time and enabled us to keep better track of intermediate results, but required multiple executions of the program with different inputs. This happened in particular for $n=7,s=2$ and the following three sets of parameters
    $p=(d,r,f,r,v,t,u)\in\Big{\{}(4,1,0,4,1,1),\quad(4,1,0,4,2,2),\quad(4,1,1,4,1,1)\Big{\}},$
    while all others parameters for $n=7,s=2$ could be computed within a few minutes each. We will explain our strategy for the case $p=(4,1,0,4,1,1)$ in detail:
    1. (1)
      First of all, we computed all extensions to $3\times 7$ matrices that could not be excluded by validity checks.
    2. (2)
      We then resumed our calculations for these $3\times 7$ matrices. For all such extensions except for
      $\begin{pmatrix}0&0&0&1&2&4&8\\ 0&0&0&0&16&32&64\\ 0&0&0&128&256&512&1024\end{pmatrix},$
      the program terminated relatively quickly. We handled these ’easy’ extensions in parallel.
    3. (3)
      For the $3\times 7$ matrix displayed above we computed the possible extensions up to the entry $(3,4)$. This means we added a partial fourth row, whose first 4 entries are already clear by the symmetry of a quaternionic matrix and computed one further element. We obtained the 10 possibilities
      $A_{0}=\begin{pmatrix}0&0&0&1&2&4&8\\ 0&0&0&0&16&32&64\\ 0&0&0&128&256&512&1024\\ 1&0&128&1&x\end{pmatrix}$
      where we had
[…]
      These cases were then again handled in parallel, where the cases for
[…]
  3. (c)
    We also applied this strategy for parameters for $n=7,s=4$, but used even more check points in between, at which we parallelized certain calculations. While it sometimes happens that matrices with high powers of 2 occurring early in the matrix are discarded quickly due to violation of rigidity constraints, most of the time, a high power of 2 occurring in the first two rows of the corresponding quaternionic matrix indicates a huge computational effort to be made.
[…]
    parameter runtime
[…]
    Table 3. Computed parameters with longest runtimes for $n=7$ and $s\geq 4$
    We were not yet able to produce results for the remaining $4$ choices of parameters
    $p\in\{(3,1,0,3,0,0),\quad(4,1,0,4,2,2),\quad(4,1,0,4,1,1),\quad(4,2,1,4,1,1)\}.$
[…]
  4. (d)
    Finally to have an overview consisting of comparable results, we include the runtimes for various cases in Table 4, where those for $n=7,s\in\{2,4\}$ are not obtained from a single run but as the sum of all partial executions, further assuming that no computations have been executed in parallel.
    Our actual implementation of Algorithm 1 was written in Python 3.12.4 and run on a MacBook Pro with Apple M4 Pro Chip and 48 GB RAM.
@@
[…]

### v2

@@ 5. Normal Quaternionic Matrices
Let $M=(M_{i,j})_{0\leq i,j\leq n-1}$ be a normal quaternionic matrix of type $\epsilon\in\{0,1\}$ with quaternionic structure S:=S(M):=(G,1,q)S:=S(M):=(G,-1,q). Set $x:=\dim(V(-1))$ if $1\not=1\in G$ and otherwise $x:=\operatorname{max}\{\dim(V(a))\mid a\in G\setminus\{1\}\}$. We write $s:=s(S)$. Then
@@
The values $(q(a_{0},a_{j})\mid j\in J)$ must be independent for else there would be a subset $\emptyset\not=I\subset J$ such that
@@
Proposition 5.1 directly produces a list of ’potential’ normal quaternionic matrices of order $1$ and $2$ and the different matrices are clearly not isomorphic as the dimension of the value sets $V(a)$ is invariant under isomorphism. For type 0, we get exactly the matrices
@@
Given $x,y\in\mathbb{N}$ we say that $y$ is a factor of $x$ if xy<xx*y<x.
@@
By Proposition 5.1, the first row of $M$ has the asserted shape, setting d:=nxd:=n-x if $s\leq 2$ and d:=nx1d:=n-x-1 otherwise.
We write S:=S(M)=(G,1,q)S:=S(M)=(G,-1,q) and fix a basis $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ of $S$ such that $M_{\mathcal{A}}(S)=M$. To keep notation brief, we usually only write $q$ for the quaternionic map instead of $q_{\mathcal{B}}$ for a basis $\mathcal{B}$ of the full 2B-group $B$ group whenever no confusion is imminent.
@@
Permuting the utu-t columns in front of the last trt-r columns of $B_{4}$, we can thus attain a block of the shape
@@
(c): By Proposition 5.1, the dimension $\dim(V(a_{0}))=n-d$ is maximal. If $d\not=n-1$, it follows that $M_{0,0}=M_{1,0}=M_{0,1}=M_{1,1}=0$ and counting the number of zeros in the second row with Proposition 5.5, we get $\dim(V(a_{1}))=n-r-v+t$. We thus have $n-r-v+t\leq n-d$ and the claim follows. ∎
@@
This can be seen as the rows with index $3$ resp. $5$ of $\overline{M}$ are (0 1 0 1 2 3 2 3)(0\ 1\ 0\ 1\ 2\ 3\ 2\ 3) and (0 2 1 3 0 2 1 3)(0\ 2\ 1\ 3\ 0\ 2\ 1\ 3), so the entries $1$ and $2$ appear in both rows but do not appear in common columns.
@@
    For direct products, the situation is more involved. Take structures S=(G,1,q)S=(G,-1,q) and $S^{\prime}=(G^{\prime},-1,q^{\prime})$ with normal quaternionic matrices $M=M_{\mathcal{A}}(S)$ and $M^{\prime}=M_{\mathcal{A}^{\prime}}(S^{\prime})$ for bases $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ of $S$ and $\mathcal{A}^{\prime}=(a_{0}^{\prime},\ldots,a_{n^{\prime}-1}^{\prime})$ of $S^{\prime}$. We conjecture that we get $M(S\times S^{\prime})=M_{\mathcal{C}}(S\times S^{\prime})$ for a basis $\mathcal{C}=(c_{0},\ldots,c_{n+n^{\prime}-1})$ that is close to a union of $\mathcal{A}$ and $\mathcal{A^{\prime}}$ in the following way: Consider $G$ and $G^{\prime}$ as subsets of $G\times G^{\prime}$ canonically. Then $c_{0}\in\{a_{0},a_{0}^{\prime},a_{0}a_{0}^{\prime}\}$ depending on the levels of $S$ and $S^{\prime}$ and
    $\{c_{1},\ldots,c_{n+n^{\prime}-1}\}\subset\{a_{i},a_{j}^{\prime}\mid 0\leq i<n,\ 0\leq j<n^{\prime}\}.$
@@
The situation for direct products also gives rise to quaternionic matrices of elementary type structures where nonzero entries appear that are no powers of 2. This is due to the fact that we have to fix 1-1 as the first basis element. For example, we get
@@ 6. The Computation
We start with the following reduction technique. Once the ETC has been confirmed for order up to n1n-1, in order to verify the ETC for order $n$, it is enough to show the ETC for nondegenerate structures that are neither rigid nor pythagorean by Proposition 2.8(a) and (d). Furthermore, depending on the level, we can also exclude some $2$-rigid cases. To make use of this, we include the following results that show how to read these properties off normal quaternionic matrices.
@@
Let $S$ be a quaternionic structure of order $n$ with normal quaternionic matrix $M$. Then $S$ is nondegenerate and neither rigid nor pythagorean if and only if every row of $\overline{M}$ has at least $4$ and at most $2^{n-1}$ zeros.
@@
The first part follows immediately from the definitions as the zeros in a row of $\overline{M}$ correspond 111-1 to the elements of $V(a)$ for some $a$. Considering Proposition 5.1 along with (b)(M2) and (b)(M3), we get the second part of the assertion. ∎
@@
Let S=(G,1,q)S=(G,-1,q) be a nondegenerate quaternionic structure of order $n$ with normal quaternionic matrix $M$. We have $s(S)=1$ if and only if $M$ has only zeros on the diagonal. Otherwise we have $s(S)=2$ if and only if $M_{0,0}=0$. Lastly, we have $s(S)=4$ if and only if $M_{0,0}=1=\overline{M}_{I,J}$ for some $I,J\subset\{0,\ldots,n-1\}$ such that $\overline{M}_{\{0\},I}=\overline{M}_{\{0\},J}=0$.
@@
$1\in D(\langle 1,1\rangle)\iff q(-1,-1)=0\iff M_{0,0}=0.$
If $s(S)=4$, we have $M_{0,0}=1$ as seen before and $1\in D(\langle 1,1,1,1\rangle)$. Note that by the correspondence of $q$ and $M$ (cf. 4.4) the existence of $I,J$ as asserted above is equivalent to the existence of $a,b\in G$ such that $q(a,b)=1$ and q(1,a)=q(1,b)=0q(-1,a)=q(-1,b)=0. By a classical property which also holds in the abstract setting (cf. [Mar80, Proposition 2.10]), we have
@@
Hence, we find $a\in D(\langle 1,1,1\rangle)$ and $b\in D(\langle 1,1\rangle)$ such that q(a,1)=q(b,a)=q(1,b)=0q(-a,-1)=q(-b,a)=q(-1,b)=0. By bimultiplicativity and symmetry of $q$, and applying q(1,1)=1q(-1,-1)=1, it follows that q(1,a)=1q(-1,a)=1 and $q(a,b)=q(a,-1)\ast q(a,-b)=1$. As pointed out above, the existence of $a,b$ proves the existence of $I,J$ as asserted. Conversely, if there are $a,b\in G$ such that $q(a,b)=1$ and q(1,a)=q(1,b)=0q(-1,a)=q(-1,b)=0, we can do the same computation as above to show that $1\in D(\langle 1,a\rangle)$ and $a\in D(\langle 1,1,1\rangle)$, completing the proof. ∎
@@
The previous two results enable us to apply the results stated in Proposition 2.8 to exclude some cases. We explicitly collect these here for a better overview. To confirm the ETC for order $n$, whenever we know that the ETC holds for order up to n1n-1, we can exclude quaternionic structures whose normal quaternionic matrix $M$ has one of the following properties:
  • There is a row of the completion $\overline{M}$ that has only zeros or fewer than 44 zeros, cf. 2.8 (a), (d) (d)(i) and (d)(ii).
  • We have $M_{n-1,n-1}=0$ (i.e. level 11) and any row of the completion has fewer than 88 zeros, cf. 2.8 (d) (d)(iii).
  • We have $M_{0,0}$ (i.e. level at most 22) and any row of the completion starting with 0 has fewer than 88 zeros, cf. 2.8 (d) (d)(iv).
  • We have $M_{0,0}=M_{1,2}=1$ and $M_{0,3}=2$ or $n=3$. This means that 1-1 is $2$-rigid and we have level $4$, cf. 2.8 (d) (d)(v). This corresponds to the parameters (cf. Proposition 5.5) d=n3d=n-3, $f=1$, $r=1$ arbitrary $v,t,u$. Note that in this case, we always have v=n2v=n-2 by the restrictions presented in Corollary 5.6.
@@
All quaternionic structures of order up to $7$ are of elementary type.
@@
By the results of Carson and Marshall as cited in Proposition 2.8(c), every quaternionic structure of order up to $5$ is of elementary type. Hence, to prove the ETC for order $6$, it suffices to compute the correct upper bound for the nonpythagorean, nonrigid and nondegenerate normal quaternionic matrices, also excluding the existence of $2$-rigid elements in the cases mentioned in Lemma 2.9. By Remark 2.10 there are $81=\sum_{k>1}e_{k}(6)$ nonrigid and nondegenerate normal quaternionic matrices of order $6$. One readily checks that among those, there are $10$ of level $1$, $28$ of level $2$ and $43$ of level at least $4$, see Appendix A.
@@
Applying this, we verified the ETC for order $6$ by running Algorithm 1 with the above restrictions, computing that there are at most $3$ resp. $19$ resp. $43$ normal quaternionic matrices of order $6$ and level $1$ resp. $2$ resp. at least $4$.
The argument for quaternionic structures of order $7$ is the same, building further on the previously stated results for order $6$. Here, the lower bound on the number of normal quaternionic matrices given by the elementary type structures further fulfilling the restrictions coming from Proposition 2.8 is $4$ resp. $65$ resp. $199$ for level $1$ resp. $2$ resp. at least $4$. Again, these numbers are obtained as upper bounds by Algorithm 1. ∎
We briefly outline some pseudo code below in Algorithm 1 to classify normal quaternion matrices while the complete version we used can be found via the following link: https://github.com/nicolo314/NormalQuaternionicMatrices.
In the program, partial quaternionic matrices (or what is aimed to be completed to such) are created row by row and via depth search in order to minimize the data that has to be stored simultaneously. To keep track of the current object, we save our current data in an $n\times n$ matrix $M$ whose entries are tuples. The first entry of the tuple in $M(i,j)$ is a partial quaternionic matrix $A$ with entries up to position $(i,j)$ and the second entry of $M(i,j)$ is a list of possible extensions at index $(i,j+1)$ of $A$ that still have to be handled. We make use of functions that increase resp. decrease $(i,j)$ to the next resp. last index, taking into considerations that elements of a quaternionic matrix up to the diagonal of a given row are already determined by previous entries.
@@
     if Validity Check(AA) = true  then
      $A\leftarrowAdd$ new row(AA)
@@
      $M(i,j)\leftarrow(A,\,$possible next elements(AA)))
@@
        else if j=n1j=n-1 then
         if i=n1i=n-1 then
           if Validity Check(AA) $\textbf{true}$ then
@@
           $A\leftarrowAdd$ new row(AA)
           $M(i,j)\leftarrow(A,\,$possible next elements(AA)))
@@
   return AA appended by partial row $k+1$ up to diagonal using (b)(M2) and (b)(M3)
function possible next elements(AA)
  if Validity Check(AA) = true  then
@@
In order to tap the full potential of a modern computer architecture with multiple cores, we parallelized some of our calculations. This is not yet necessary for the case of order $6$ which was computed in about 14 seconds without applying any parallelization. However, dealing with the case of order $7$, the runtimes get tremendously large as can be expected due to the exponential growth both of the number of cases to be considered and the size of the completion of the quaternionic matrices.
To have an overview consisting of comparable results, we include the runtimes for various cases in Table 3, where those for $n=7,s\in\{2,4\}$ are not obtained from a single run but as the sum of all partial executions, further assuming that no computations have been executed in parallel.
Our actual implementation of Algorithm 1 was written in Python 3.12.4. The cases in which $n\leq 6$ or $s<2$ as well as most of the parameters with $n=7$ and $s=4$ were run on a MacBook Pro with Apple M4 Pro Chip and 48 GB RAM. The most time consuming parameters were computed using the HPC cluster Elysium of the Ruhr-Universität Bochum.
@@
2 $\units0\,\unit${s} $\units0\,\unit${s} $\units2\,\unit${s} 33 days $\unith15\,\unit${h}
4 $\units0\,\unit${s} $\units0\,\unit${s} $\units12\,\unit${s} $204\,\text{days}$